OpenStudy (anonymous):

The circumference of a circle is 8pi inches. find its radius and area.

OpenStudy (anonymous):

Hint: \[\Large circumference=2\pi(radius)\\ \Large Area=\pi r^{2}\]

OpenStudy (anonymous):

can you walk me through it

OpenStudy (anonymous):

you have the circumference. Plug it into the equation above & simplify to get the radius. Once you have the radius, compute the area using the second equation.

OpenStudy (unklerhaukus):

\[C=2\pi r\] \[\frac{C}{2\pi}=r\]

OpenStudy (anonymous):

unkleRhaukus what is the area

OpenStudy (anonymous):

c/2pi is not correct

OpenStudy (anonymous):

It most certainly is.

OpenStudy (anonymous):

it should be correct

OpenStudy (anonymous):

i mean, c/2pi gives the radius in inches.

OpenStudy (unklerhaukus):

\[A=\pi r^2\]

OpenStudy (unklerhaukus):

\[A=\pi\left(\frac C{2\pi}\right)^2\]

OpenStudy (anonymous):

look if the circumference in this case is 8 and the circumference of a circle is \[2\pi r\] then in this case \[2\pi r =8\pi\] so \[r = 8\pi/2\pi \] then put the value of r in the area equation

OpenStudy (anonymous):

I do not know what is going on but i am going to do another similar problem because they say both answers are wrong

OpenStudy (anonymous):

hello you still there

OpenStudy (unklerhaukus):

.

OpenStudy (anonymous):

circumference of a circle is 8pi inches: C = 2π r = 8π -> r = ....? @cdelomas Can you find the radius?